Barney decides to quit his job as a corporate accountant, which pays $12,000 a month, and goes into business for himself as a certified public accountant.

He runs his business from his converted garage apartment, which he could rent out for $325 a month if he wasn’t using it as a home office. He must purchase office supplies worth $80 a month, and his monthly electricity bill has increased by $60 now that he is working out of his home office.

After six months of working from home, Barney has earned an average of $15,000 per month.

Instructions: Enter your answers as a whole number.

a. What are Barney’s monthly explicit costs?

     $   

b. What are Barney’s monthly implicit costs?

     $   

c. What are Barney’s monthly economic costs?

     $   
Answer #1

a) Barney's monthly explicit costs= 325+80+60= 465

b) Barney's monthly implicit costs= salary earned by the previous jobs= 12000

(Implicit costs are costs of a possible benefit that did not occur)

c) Barney's monthly economic costs= explicit costs+implicit costs= 465+12000= 12465